OpenStudy (anonymous):

Graph the equation by plotting 3 points -5x - 3y=-7

OpenStudy (anonymous):

use y=0 y=1 y=2 and solve the equation for x from there

OpenStudy (anonymous):

y=o x=7/5. x=0 y=7/3 x=1 y=2/3

OpenStudy (anonymous):

wait that can't be right

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

y=2/3 <--- it has to be a whole number not fractions

OpenStudy (anonymous):

because on my graph it wont show up

OpenStudy (anonymous):

what do i multiply?

OpenStudy (anonymous):

does any one know how to do this?

OpenStudy (anonymous):

Nothing. Are you sure you've posted the correect formula? Because when x=1 y definitely = 2/3

OpenStudy (anonymous):

thats how its written